The pressure in a constant-volume gas thermometer is 0.700 atm at 100°C and 0.512 atm at 0°C.

(a) What is the temperature when the pressure is 0.0800 atm?
°C

(b) What is the pressure at 670°C?
atm
Answer #1

PF 6.700 atm TE IUC = lo6+273 k = 373 k P = 0.512 at 7, = 6c +273 le T2= 273 k. =) PV=nRT : Ij constant & Pr TIT cu ( B= 6.